Reports of 21032
FCL (Air)
shinji.miyoki - 21:28 Saturday 26 November 2022 (23047) Print this report
EYV temp adjust

Because EY F1 GAS seems to saturate, I reduced the heater power from 600W to zero in EYV.

Because 600W is too much, we need to add a heater to adjust  0~500W range.

MIF (ITF Control)
shinji.miyoki - 19:12 Saturday 26 November 2022 (23046) Print this report
Comment to Recovery of FPMI (23041)

The monkey will climb up?

IOO (OMC)
kenta.tanaka - 14:10 Saturday 26 November 2022 (23045) Print this report
OMC Lock after OMC chamber evacuation and OMC REFL WFS finalization

Yokozawa, Tanaka

### Abstract

I locked OMC after vacuuming the OMC chamber. I measured the open-loop transfer function of the OMC LSC loop. The UGF is tuned around 100 Hz.

Now the transmitted power of OMC is 45 mW, while IMC's transmitted power is 2.2 W and X-arm is locked. According to previous Micheal-san's estimation, TEM00's power is 40 mW while IMC's transmitted power is 2.9 W, so the transmitted power of OMC is increased even though IMC's transmitted power is decreased. Roughly speaking, 63 mW at 2.9 W is 47 mW at 2.2 W, and if 45 mW is transmitted, the transmittance is 96%.

I tried to measure the finesse of OMC with the ringdown method to check whether the mirrors were contaminated during the vacuuming process. However, Yokozawa-san pointed out the required decay time is too short (~1us, ref. klog) to measure with DGS. So I gave up measuring the finesse with the ringdown method by DGS.

We finalized the OMC REFL path and did phasing about both OMC REFL WFSs

### What I did

#### Lock OMC and measure OLTF

  • After vacuuming the OMC chamber, I aligned the downstream from SR2 while X-arm was locked. Thanks to Ushiba-san, I can use OMMT2 TRANS QPDs to align OMMT1. So, I centered the beam on OMMT2 TRANS QPDs with OMMT1. After then, I centered OMC QPDs with OMMT2 and OSTM.
  • I engaged the loop from OMC QPDs to OMMT2 and OSTM to fix the beam position on their QPDs.
  • I scanned the length of the OMC cavity with HV PZT1 monitoring the transmitted beam shape on the OMC TRANS camera. TEM00 is flashed, and the trigger locks OMC. At first, PZT seemed to saturate. so I tuned the gain not to saturate PZT.
  • After the gain tuning, we measured the open-loop transfer function to see if we could include a boost filter. We confirmed a phase margin even with the boost filter, so we added the boost filter and measured the open-loop transfer function again(Fig. 1). Finally, the UGF is around 100 Hz, and the phase margin at 100 Hz is around 40 degrees.
  •  In this condition, the alignment was adjusted by changing the set point of the QPD loop to maximize the amount of transmitted light. Finally, the transmitted light of the OMC (K1:OMC-TRANS_DC_SUM_OUT_DQ) was 45 mW. (Fig. 2) According to previous Micheal-san's klog, TEM00's power is 40 mW while IMC's transmitted power is 2.9 W, so the transmitted power of OMC is increased even though IMC's transmitted power is decreased. Roughly speaking, 63 mW at 2.9 W is 47 mW at 2.2 W, and if 45 mW is transmitted, the transmittance is 96%.

#### Finesse measurement

  • I tried to measure the finesse of OMC with the ringdown method to check whether the mirrors were contaminated during the vacuuming process.
  • I closed the PSL shutter while OMC was locked and measured the ringdown of OMC's transmitted power.
  • Fig. 3 shows the result. there is a strange step in the ringdown shape of transmitted power, and the decay time estimated from the measured ringdown is several ten seconds. It seems too long. And Yokozawa-san pointed out the required decay time is too short (~1us, ref. klog20764) to measure with DGS. So I gave up measuring the finesse with the ringdown method by DGS.

#### OMC REFL WFS finalization

  • After the OMC lock, We took fine alignment of the OMC REFL path.
  • I turned on PZT drivers for PZT DC loops. After then, I tweaked the mirrors to center both OMC REFL WFSs roughly.
  • I injected the reflection beam into OMC REFL camera.
  • I did the phasing about both OMC REFL WFSs. Fig. 4 shows the results.
Images attached to this report
PEM (Center)
takaaki.yokozawa - 8:31 Saturday 26 November 2022 (23044) Print this report
Microphone test
For testing the performance of microphone, we lent several microphones, during this test, we couldn't use following microphone (1 week?)

K1:PEM-MIC_PR_TABLE_ISS_Z_OUT
K1:PEM-MIC_PR_BOOTH_REFL_Z_OUT
K1:PEM-MIC_BS_FIELD_BS_Z_OUT
MIF (ITF Control)
tomotada.akutsu - 1:02 Saturday 26 November 2022 (23043) Print this report
Comment to Recovery of FPMI (23041)

Cool UI!

MIF (General)
takafumi.ushiba - 23:41 Friday 25 November 2022 (23042) Print this report
PRMI recovery Day2

Continued from klog23023.

Abstract:
I continued the recovery of PRMI.
PRMI 3F was recovered but AS port was not so dark.
I'm not sure the exact reason.
One possible reason is PRMI was locked to the bright condition at AS due to the sign mistakes.

Detail:
First, I checked the sign of PRC again after FG replacement (klog23038) but still the sign is flipped from the one on Nov. 3.
So, I implemented the new sign to the filter bank and now PRC can be locked on resonance of carrier by guardian.

After that, I locked PRMI with 1F signal and engaged the ADS but sideband build up was reaching about 0.4 this time while it was about 1 on Nov. 2 (klog22764).
Since, I faced the sign flip of PRC, so I checked the phase of I and Q phase signals, and they seems not optimized.
So, I rotated the phase from 11 degs to 71 degs to minimize the coherence between I and Q phase signals when PRMI is locked.
After that, I checked the coherence when rotating 45 degs from that position and ci\onfirmed the coherence becomes high.
Then, sideband build up (POP 90) can reach 1.

After the alignment of PRMI, I locked PRMI with 3F signals and it can be locked after tweeking the gains of MICH and PRCL loop.
However, after the lock, AS and POP becomes bright (power build up at AS is larger than at POP, fig1).
Since the sign and phase was changed after the IMC troubles, it is better to check the phase one by one. 

Note:
Figure 2 shows the good alignment for PRMI today.

Images attached to this report
MIF (ITF Control)
takafumi.ushiba - 22:32 Friday 25 November 2022 (23041) Print this report
Recovery of FPMI

I checked if FPMI can be locked after the replacement of function generator for local oscillators (klog23038).
After the usual initial alignment, I requested Observation and FPMI can be locked without any problem.

Non-image files attached to this report
Comments to this report:
tomotada.akutsu - 1:02 Saturday 26 November 2022 (23043) Print this report

Cool UI!

shinji.miyoki - 19:12 Saturday 26 November 2022 (23046) Print this report

The monkey will climb up?

IOO (General)
shinji.miyoki - 20:48 Friday 25 November 2022 (23038) Print this report
Comment to IMC cannot lock (23008)

I swapped the SG. So now N5181B for 5.6 MHz:13.3dBm, and E8663D for 13.78MHz:7.4dBm.

Both SGs are now controllable on the web. IP information is restricted only to Ushiba-kun and Yamamoto-kun, and Oshino-kun.

VIS (General)
satoru.ikeda - 18:58 Friday 25 November 2022 (23036) Print this report
Comment to Routine confirmation of DGS maintenance day (17831)

DGS Regular maintenance day(11/18)

Stepper motor
 [Checked]
  Check if the power supply can be controlled by BIO
  Check if the script can be started
  Checked that the limit switches installed are green.
 [Check result]
  No problem

Release Sensor
 [Checked the Release Sensor.]
 Checked that Release Sensor is 0, 1: Seated.
 [Check result]
 ETMX, ETMY, ITMX, ITMY : No problem

Picomotor
 [Confirmation]
  PingTest is performed.
 [Result]
  Only known

 NG:
  SR2(IM) : Not restored today due to other work priorities

  MCI,STM1,
  POP_POM,POS_POM,POM1
  MCF(MCO,MCI)
   => Confirmation only, no recovery operation was performed.

 OK:
  ETMX,ETMY,ITMX,ITMY
  BS(IM,BF),SRM(IM,BF),SR2(IM,BF),SR3(IM,BF,STM)
  PRM(IM,BF),PR2(IM,BF),PR3(IM,BF)
  MCE,IMMT1,IMMT2,OMMT1,OMMT2,OSTM
  REFL_WFS,POP,POP2,POS,POS2,POP_WFS,AS_WFS
  PCAL(EX1,EX2,EY1,EY2):Not checked because it is ON only when used

Temperature controller
 [Checked]
 Is the temperature acquisition working properly?

 [Check result] 
  Is the temperature acquisition working properly?
  => No problem

Precision air Processer
 [Check Contents]
 Is the temperature acquisition working properly?

 [Result]
  No problem

FunctionGenerator
 (Confirmation)
 Is there a ping response?
 Are all EPICS channels alive?
  [sitemap]-[Commissioning]-[ALS FG]-[Xarm],[Yarm].
 [Check result]
 X and Y PLLs are working properly.

HWP control PC
 (Confirmation)
 Check if it is possible to login to the control PC.
 (Result) 
 Both k1hwp0, 1 and 2 are OK.
 

Images attached to this comment
Non-image files attached to this comment
DGS (General)
takahiro.yamamoto - 17:50 Friday 25 November 2022 (23040) Print this report
Model rate was fixed on k1visbst and k1vissr3p
[Ikeda, Ushiba, YamaT]

Model rate of BS tower and SR3 payload models was modified because they run as a different sampling rate from other suspensions.
- BS tower: 4096Hz => 2048Hz
- SR3 payload: 2048Hz => 16384Hz

-----
Models themselves were changed just only 'rate' on the 'cdsParameters' block.
Along with these changes, the list of IPC (/opt/rtcds/kamioka/k1/chans/ipc/K1.ipc) and foton filter fines (/opt/rtcds/kamioka/k1/chans/K1VIS*.txt) were also modified.

As far as I know, the IPC list isn't modified and removed in automatically though a new connection is automatically added.
Sampling rate of IPC sender model is written in the IPC list with 'ipcRate' as a keyword. So we modified it in manually for the related IPC blocks.

Foton file can be converted to a new sampling rate by using /opt/rtcds/userapps/release/cds/common/scripts/updateRate.py.
But some information such as ramp time doesn't seem to be kept in a modified file though filter zero/pole design is kept.
For this reason, Ushiba-kun restored the missing information one-by-one in manually.
updateRate.py should be modified to keep all information. But who has a enough time...?
DGS (General)
satoru.ikeda - 16:58 Friday 25 November 2022 (23039) Print this report
Comment to k1ctr4 was stopped for a test of a new workstation (22953)

Workstations up to k1ctr0-4 in the control room have been replaced with new PCs.
k1ctr5 has flicker when connected to a monitor, so it will be replaced after the monitor is purchased.
 

Non-image files attached to this comment
LAS (High Power Laser R&D)
osamu.miyakawa - 16:39 Friday 25 November 2022 (23037) Print this report
HPL works

Uehara, Miyakawa

Measurement of IMC OLTF last week showed that a big phase delay exists on wide band EOM with the new HPL. We tried to recover the phase using short cable between RF high voltage amplifier and input of wideband EOM at the box of new HPL. We exchanged the cable length 1.5m to 0.5m but we could not see any significant phase recovery. It means that we have to modify the servo circuit.

 

We measured the reflectivity of PMC when it was locked to the new HPL. It was a huge number of ~83% by comparing the reflectivity when PMC was not locked.

Then we measured the beam profile at several points (after 1st lens, and 2nd lens) several times to match the mode. We could not have time to tune the finel lens or alignment, but the reflectivity got lower than 50%. We will tune them next Friday.

DGS (General)
satoru.ikeda - 15:55 Friday 25 November 2022 (23035) Print this report
Updating model files

Request from Toyama-u Sako-san
 Install and restart k1psliss

Request from Oshino-san
 Add DAQ channel to k1tmsy model
  K1:TMS-Y_IR_QPDA1_SUM_OUT
  K1:TMS-Y_IR_QPDA2_SUM_OUT
 k1tmsx,k1tmsy:
  Delete DAQ channel in K1:TMS-{X,Y}_{GR,IR}_QPDA{1,2}_SEG{1,2,3,4}_IN1

Request from Yamamoto-san and Ushiba-san
 1. update, install and restart k1aso and k1asc
  Addition of OMMT2 TRANS QPD1 and QPD2

 2. model rate change
 k1visbst: change rate from 4K to 2K
 k1vissr3p: change rate from 2K to 16K
 /opt/rtcds/kamioka/k1/chans/ipc
  => Manually rewrite the rate of the channel to be sent from k1visbst  
 After this, k1visbsp was also re-built and restarted.
 

Images attached to this report
Non-image files attached to this report
VIS (EY)
lucia.trozzo - 15:42 Friday 25 November 2022 (23034) Print this report
Transfer function from IP L to TM L

This klog refers to the past days' activity.
I measured the Tf from IP_L to TM_L, injecting white noise in the frequency range [0 10] Hz while the IP was under Inertial control.
Pic 1 shows the result of the measurement. Here it is possible to see that, below 1.5 Hz where the mechanical resonances of the system are well-known, the coherence is almost 1.
To project the IP contribution on Darm, I integrated the low-frequency fit (below 1.5 Hz) whit the simulated model calculated with the Model plotter.
Pic2 shows the comparison of the TF and the model I made.
Pic3 shows the projection of the IP motion on Darm, calculated following this klog.

Note: this is just a preliminary result.
The diaggui file of the TF is in the following directory:
kagra/Dropbox/Subsystem/VIS/vis_commissioning/IP_diag/TF_IP2TMCL

Images attached to this report
VIS (EX)
lucia.trozzo - 14:34 Friday 25 November 2022 (23033) Print this report
Comment to New IP blending with the Geophones and the LVDTs sensor corrected has been tested and implemented (22873)

[Takahashi, Lucia]

We reshaped the blending filters for the L and T directions to reduce a bit more the re-injection of the seismic noise in the region [0.4, 1] (See Pic1, Pic2, Pic3). We implemented these new filters and we measured the spectra. Pic4 shows the H4 spectrum when LP80_new and LP80_rev are running. It seems to be working fine.
Pic5 and Pic6 show the BF spectra along L and T directions.
Pic12, Pic13, and Pic14 show the TM spectra along L, P, and Y directions. No evidence of re-injection of geophone noise below 0.1 Hz has been detected.
Note:
We have already implemented and tested this new blending strategy for L and T in all the TypeA suspensions.
In the MEDM screen, those filters are saved in the filter bank FM6 and called LP80_rev and HP80_rev.
The previous strategy is saved in the filter bank FM5 and the filters are saved as LP80_new and HP80_new.

Images attached to this comment
IOO (General)
shinji.miyoki - 9:50 Friday 25 November 2022 (23032) Print this report
Comment to IMC cannot lock (23008)

Amplitude measurement for the Wenzel harmonics generator was done as reported, while the phase measurement was not done. High multiple harmonics could have phase delay changes according to their amplitude.

VAC (BS)
nobuhiro.kimura - 8:50 Friday 25 November 2022 (23031) Print this report
Completed of vacuum leak test for a new ion pump near GVIY-BS

When: 22/Nov. /2022
Who: Kimura, Yasui and M. Takahashi
Task:
We performed a vacuum leak test  for a new ion pump near GVIY-BS.

There were no leakage<1 x 10^-13 Pam^3/s.

After the leak test, we opened GV between the vacuum duct and vacuum pump unit. 

 

Images attached to this report
CAL (XPcal)
dan.chen - 8:15 Friday 25 November 2022 (23030) Print this report
Adjustment for Pcal beam power at ETMX

Date: 2022/11/25 early morning

Summary:

- Optical efficiencies of XPcal were updated.
    - The overall optical efficiency is 0.966, so that I set 0.983 into e_Tx and e_Rx of path 1.
    - The overall optical efficiency is 0.960, so that I set 0.980 into e_Tx and e_Rx of path 2.
- The gain of EX_2_INJ_V was adjusted to have the same power at ETMX.
    - The value is 0.961.
- The separation rations were also updated from 0.5 to 0.501 for path 1 and 0.499 for path 2.

A detailed report will be here.

MIF (ASC)
hirose.chiaki - 8:05 Friday 25 November 2022 (23029) Print this report
Comment to Installed ASC of TMSY DCQPDs ->Yarm : It is for initial alignment of measure WFS offset (22972)

The picture describing the process for passing LOOP is when it is done before turning on LOOP. It is not the process after it is used. My English was wrong.
We will also restore it after the measurement is completed.(I'm sorry for any inconvenience this may cause when inithal alignment.)

VAC (General)
shinji.miyoki - 7:59 Friday 25 November 2022 (23024) Print this report
Another leak in OMM tank

After today's Kimura-san's vacuum leak repair in the OMC and OMMT tank, TMP was successfully turned on (This report should be done by Kimura-san).

However, the vacuum level stopped at 0.12 Pa suddenly and kept constant. 

I re-entered the KAGRA center area and I confirmed that the TMP/Scroll pumps were working. So actually other leaks should exist.

I tried the "ethanol method", and found a leak at the GV-side d400 flange on the OMMT tank because the vacuum level enhanced from 0.12 Pa to 0.6 Pa when I put ethanol. Also I found several suspicious blank flanges and a connector flange on the OMMT and OMC tank.

I hope they will be repaired by Kimura-san.

Comments to this report:
nobuhiro.kimura - 7:48 Friday 25 November 2022 (23027) Print this report

When: 24/Nov./2022
Who: Kimura, Yasui and M. Takahashi
Task:

 We identified two leak points by spraying alcohol.  The leaks were in the K-400 flange joints with metal gaskets.

We retightened the flange joints at three locations, including these two, to stop the leaks.

For each K-400flange, 16 clamps were added and tightened to a torque of 60 Nm.

The remaining one flange was an ICF203 flange and was retightened with a torque of 15 Nm.

MIF (ASC)
hirose.chiaki - 7:57 Friday 25 November 2022 (23026) Print this report
Measurement of stability to sensing matrix of Yarm

Tanaka, Hirose

We have measured stability of sensing matrix of Yarm since 20th Nov. (I'm so sorry for missing to post klog.)
The purpose is to research the effects of birefringence by measuring the stability of the sensing matrices due to changes in the beam spot of ITMY.
How to measure is following.

  1. Yarm is turn to IR lock.
  2. Loop (DCQPDs of TMSY ->IY, EY) is turned on.
  3. Record and offroad good oplev of IY, EY.
    → To adjust the optical axis to the same optical axis each time to start measurement.
  4. Loop (DCQPDs of TMSY ->IY, EY) is turned off and OLDAMP (from TM to MN) is turned on. (For detail is klog22972)
    → Now TMSY returns more noise than OLDAMP because it is not covered. After setting GOOD OPLEV, switch to OLDAMP control.
  5. ADS(TMSY->IMMT2,PR2) is turned on. Wait for ADS to settle.
  6. We measure the sensing matrix with excitation applied to each degree of freedom by diaggui at frequency higher than the ADS control band.
    → Once the incident optical axis and the resonator axis are aligned, we measure the sensing matrix centered on that position.
  7. Change the beam position by changing the oprev set point, then measure like step6.

We will post the results later.

Images attached to this report
IOO (General)
shinji.miyoki - 7:49 Friday 25 November 2022 (23028) Print this report
Comment to IMC cannot lock (23008)

According to this klog#8247, a proper input amplitude value for the Wenzel harmonics generator is desirable to obtain enough amplitude for some harmonics.

Then klog#8422 was tried (detail measurement condition) and 13.3dBm seemed to be proper to obtain "specified values by Wenzel Inc." for some harmonics.

The data that was provided by Wenzel Inc. is here.

VAC (General)
nobuhiro.kimura - 7:48 Friday 25 November 2022 (23027) Print this report
Comment to Another leak in OMM tank (23024)

When: 24/Nov./2022
Who: Kimura, Yasui and M. Takahashi
Task:

 We identified two leak points by spraying alcohol.  The leaks were in the K-400 flange joints with metal gaskets.

We retightened the flange joints at three locations, including these two, to stop the leaks.

For each K-400flange, 16 clamps were added and tightened to a torque of 60 Nm.

The remaining one flange was an ICF203 flange and was retightened with a torque of 15 Nm.

IOO (General)
shinji.miyoki - 23:49 Thursday 24 November 2022 (23025) Print this report
Comment to IMC cannot lock (23008)

Not 13.3 dBm but 10.0 dBm is selected for 5.6MHz signal.

 

MIF (ITF Control)
takafumi.ushiba - 22:36 Thursday 24 November 2022 (23023) Print this report
PRX/Y and PRMI recovery

Abstract:
I checked PRX and PRMI can be locked or not by guardian.
PRX can be locked from guardian with both 3F and 1F.
Also, PRMI with 1F can be locked by guardian but PRMI with 3F cannot be locked.
After several investigation, I found that sign of REFL135 was somehow flipped.
I'm not sure why this happens and it is necessary to investigate.

Detail:
I checked MICH, PRC and PRMI can be locked by guardian, which could be on Nov. 3.
MICH with 3F, PRX with 3F, PRX with 1F, and PRMI with 1F can be locked but PRMI with 3F cannot (I forgot to check MICH with 1F but probbly can).

First, I thought it could be alignment, so I engaged ADS to IMMT1, IMMT2, and PRM while locking PRMI with 1F.
After the ADS, sideband bildup (POP90) became 0.2 from 0.05 (fig1).
Although POP 90 signals was about 1 on Nov. 3, I don't mind the absolute value because current LO power (10 dBm) was smaller than nominl (13.3 dBm).

After the alignment, Itried to lock PRMI with 3F signals but couldn't.
So, I checked MICH and PRCL one by one.
I found no strange behaviour on MICH 3F lock, however, I found PRX with 3F signals was somehow locked to anti-resonance of carrire light, which should be resonance for the carrire.
I'm not sure why this sign flip was occured (can LO change flip the sign of REFL135 signals ?).
Anywa, I fliped the sign of calibration factor for REFL135 and now PRX can be locked on resonance for the carrire by guardian.

We will continue PRMI recovery tomorrow.

Images attached to this report
Search Help
×

Warning

×